TV advertising agencies face increasing challenges in reaching audience members because viewing TV programs via digital streaHelp with b) please. It is not 462 nor 460
Answer #1

Solution :

\hat p = 0.52

1 - \hat p = 0.48

margin of error = E = 0.06 / 2 = 0.03

Z\alpha/2 = 2.576

sample size = n = (Z\alpha / 2 / E)2 * \hat p * (1 - \hat p )

= (2.576 / 0.03)2 * 0.52 * 0.48

= 1841

sample size = n = 1841